OpenStudy (anonymous):

Vehicle Stopping Distance Taking into account reaction time, the distance d(in feet) that a car requires to come to a complete stop while traveling rmiles per hour is given by the function d(r) = 6.97r – 90.39 (a) Express the speed r at which the car is traveling as a function of the distance d required to come to a complete stop. (b) Predict the speed that a car was traveling if the distance required to stop was 300 feet. could someone tell me the steps of how to go about this

OpenStudy (anonymous):

lol i thaught d(r) = 6.97r – 90.39 is the answer to part a XD

OpenStudy (anonymous):

i have no idea. i don't even know where to start.

OpenStudy (anonymous):

do you know how much marks is this question?

OpenStudy (anonymous):

a good amount im sure. i think it's asking me to find r(d) maybe?

OpenStudy (anonymous):

i can do that but it'll be like 1 mark haha

OpenStudy (anonymous):

r(d) = (d +90.39)/6.97 all i did was rearrange the formulae

OpenStudy (anonymous):

but is does help for part be so i guess we might be on the right track xD

OpenStudy (anonymous):

r(300) = (300 + 90.39)/6.97 r(300) = 390.39/6.97 r(300) = 56.01mil per hr

OpenStudy (anonymous):

that would be my answer to the question :)